OpenStudy (anonymous):

Miss Muffet always likes to have things "just so". In fact, the only way she will consume her curds and whey is 2 Units of Curds with 1 Unit of Whey. I know the Utility function is U=min(2W, C) Miss Muffet has $12. P for curds is $1, and P for Whey is $2. What is the utility maximizing bundle?? and how!?

OpenStudy (anonymous):

\[Pc \times C + Pw \times W = M \] so \[C +2W=12\] and also 2W=C so solving above 2 equations, v get 4W=12 W=3 n C=6